# Beyond the Visible: A Survey on Cross-spectral Face Recognition

### v3

@@ Abstract
{highlights}
Formalization of the cross-spectral face recognition problem
Appropriate electromagnetic spectral band for related cross-spectral face recognition
General components of a cross-spectral face recognition system
Recent advances based on deep learning in addressing cross-spectral face recognition
Future research for addressing challenges related to cross-spectral face recognition
@@ 1 Introduction
Face Recognition (FR) has been a very active research field for the last several decades [1, 2, 3, 4]. Advances in deep convolutional neural networks (CNNs) and the associated seminal work on DeepFace [5] have brought significant progress in this area, tackling a number of challenges including variations in pose, illumination and expression (PIE), as well as resolution and unconstrained settings. While such work have predominantly focused on the visible (VIS) spectrum, considering additional spectra allows for increased robustness [6, 7], in particular in the presence of different poses, illumination variations, noise, as well as occlusions. Further benefits include incorporating the absolute size of objects, as well as robustness to presentation attacks such as the use of makeup and masks to circumvent a system [8]. Therefore, comparing RGB face images against face images acquired beyond the visible spectrum, often referred to as Cross-spectral Face Recognition (CFR), is of particular significance in designing FR systems for defense, surveillance, and public safety [6] and falls under the broader category of Heterogeneous Face Recognition (HFR) that addresses the broader concept of recognizing faces across diverse and non-uniform conditions. Specifically, heterogeneous indicates that face images emerge from different sources and sensing modalities, e.g., visible, infrared, 2D-3D, resolutions, and hand-drawn sketches (see Figure 1).
The focus of this survey is on CFR, with a specific emphasis on comparing visible face images with images acquired beyond the visible spectrum, viz., the infrared spectrum. We note that CFR is more challenging than traditional FR for both human examiners as well as computer vision algorithms. This is due to following three factors. Firstly, large intra-spectral variation, where within the same modality, face samples of the same subject can exhibit larger appearance variations than face samples of different subjects. Secondly, the modality gap is of concern, where appearance variation between two face samples of the same subject can be larger than that of two samples belonging to two different subjects (see Figure 1). This can result in degraded face recognition performance. Finally, the limited availability of training samples of cross-modality face image pairs can make it difficult to design effective CFR methods using deep neural networks.
Figure 1: Examples of heterogeneous face images of the same subject captured in different modalities, i.e., domains. Images are from the Tufts dataset [9].
@@
Recent advances in CNNs and generative adversarial networks (GANs) have allowed for remarkable improvements in CFR [10, 11, 12, 13]. The goal of this survey article is to provide an overview of these advances. Specifically, we describe the proposed methods, used datasets, as well as the results and associated insights. In this context, we consider the electromagnetic spectrum ranging from near infrared (NIR) to more challenging bands such as midwave infrared (MWIR) and longwave infrared (LWIR), as illustrated in Figure 2. Unlike other introductory overview articles including [14, 6, 15], we focus on highlighting recent advances based on deep learning in addressing CFR, and techniques that have been developed to bridge the modality gap.
@@
The remainder of this paper is organized as follows. We formalize CFR in Section 2 and present some applications in Section 3. We discuss the various spectral bands, i.e., sensing modalities, that are relevant to face recognition, in Section 4. In Section 5 we describe the general modules of a CFR system, viz., face image pre-processing and feature extraction. The various deep learning based CFR methods that have been proposed in the literature are explained in Section 6 and 7, while the relevant datasets used for research in this field are presented in Section 8. Finally, in Section 9, we discuss open research problems that are currently being addressed in the field of CFR.
@@ 3 Applications of CFR
FR-systems have been widely deployed and operate mainly in the visible spectrum [19], due to the ubiquity of low-cost cameras. As motivated in the introduction, CFR-systems have been beneficial in additional applications, for example, in Intelligence Video Surveillance (IVS). Beyond access control, the application of CFR has strong potential to provide new capabilities in law enforcement, military and the AI community, especially for commercial purposes.
A pertinent application for CFR has been in law enforcement and homeland security [20]. In this context, IR-sensors can be used for long-range monitoring and capturing biometric modalities at variable standoff distances, as well as in low-light and night-time scenarios. In such cases, the acquired image often has to be compared against a gallery consisting of visible spectrum face images. Similar applications are found in military settings involving identifying individuals in challenging environments [7]. On a wider scale, HFR (which is a superset of CFR) is beneficial for deflecting presentation attacks [21, 22] or for sketch-visual identification, e.g., finding an individual whose appearance was sketched and whose face image may be present in a mugshot suspect database [23].
@@ 5.3 Face Comparison
Algorithms of different Spectra
@@
6 Reflective IR-to-Visible FR
6.1 Near Infrared
@@
IDNet. Reale et al. [45] utilized GoogLeNet networks with small convolutional filters, pre-trained on the visible CASIA WebFace dataset. For NIR-to-VIS CFR scenario, two identical networks were initialized based on the pre-trained network by excluding fully connected and softmax layers. The outputs from these two networks, i.e., VisNet and NIRNet, were concatenated to create a siamese network with contrastive loss. Such a coupled deep network design was able to map NIR and VIS faces into a spectrum-independent feature space.
@@
CFC. He et al. [12] proposed a GAN, termed Cross-spectral Face Completion (CFC), to synthesize VIS images from NIR images. The heterogeneous face synthesis process was divided into two complementary components: (a) a texture inpainting component that aims to recover a VIS facial texture map from a given NIR image; and (b) a pose correction component that aims to map any pose in NIR images to the frontal pose in VIS images. A warping procedure was used to fuse these two components in the adversarial network, which was regularized by the combinations of UV loss, adversarial loss and L1 reconstruction loss. In summary, the main contributions of this work had to do with creating a novel GAN-based end-to-end deep framework for cross spectral face synthesis without assembling multiple image patches. It contained encoder-decoder structured generators and two novel discriminators to fully consider variations of NIR and VIS images. This was the first time that unsupervised heterogeneous face synthesis problem was simplified to a one-to-one image translation problem. The decomposition of texture inpainting and pose correction enables the generation of realistic identity preserving VIS face images possible.
@@ 6.2 Shortwave Infrared
VGGFace. Bihn et al. [51] examined the ability of using pretrained VGG-Face network on visible images to extract features on SWIR images. Hypothesized was that the VGG-Face network would perform well on face images taken from both, VIS and SWIR wavelengths since their facial appearance difference is not as significant as these images captured at MWIR and LWIR wavelengths. SWIR wavelengths at 935 nm, 1060 nm, 1300 nm, 1550 nm were evaluated, as well as a composite image formed by combining 1060 nm, 1300 nm, and 1550 nm. Deep features extracted from the fc7 layer of VGGFace network were used, resulting in a 4096-dimensional feature vector. Euclidean distance calculated between feature representations of VIS and SWIR images was used as a similarity comparison.
@@
7 Emissive IR-to-Visible FR
7.1 Midwave Infrared
Appearance difference between MWIR and VIS is more pronounced than NIR and SWIR (Figure 4). Therefore, learning a common feature subspace for MWIR-to-VIS is a challenging problem. In most cases, methods developed over the years have relied on GANs to synthesize visible face images from thermal face images. Illustrative examples are shown in Figure 13 and 15. We report the related MWIR-to-VIS face comparison performance in Table 6.
@@ 7.2 Longwave Infrared
DCA. Kantarcı et al. [61] presented a Deep Convolutionnal Autoencoder (DCA), specifically based on the U-Net [58] architecture, including Up Convolution and Difference of Gaussians (DoG), which was modified it for the purpose to learn the non-linear mapping between visible and thermal face images for CFR. In the meantime, they showed that applying preprocessing and alignment would help to close the gap between these two domains and was able to improve the performance further. The proposed approach was extensively tested on three publicly available thermal-visible datasets: Carl [109], UND-X1 [108] and EURECOM [114]. In addition, authors have manually annotated $6$ facial landmarks on the Carl and EURECOM datasets in order to align the faces and assess alignment’s impact on the performance (we observe that alignment increased the performance by around $2\%$). They conduct experiments with three different setups: the first two experiments investigated the effect of the decoder method, DoG filter were not applied to images as preprocessing step. For the third experiment, the effect of the DoG filter was tested on the Up Convolutionnal Autoencoder Model, proposed in this work.
@@
DAF. Fondje et al. [118] proposed a Domain Adaptation Framework (DAF) with a new feature mapping sub-network including (a) a cross-domain identification-loss function for effective “co-registration” and “synchronization” and (b) a domain invariance-loss function for cross-domain regularization. Within each domain, the feature representations were acquired using a truncated pre-trained CNN (based on VGG16 or ResNet50). The authors were interested in finding the optimal depth-level of the network while preserving the meaningful discriminative information shared by both the visible and thermal spectra. DAF is then embedded in their proposed Residual Spectral Transform (RST), which is a residual block. The key role of RST is to transform the features with the help of three particular 1x1 convolutions while enhancing the discriminability. Experiments were conducted using three datasets (considering only thermal probes): the polarimetric thermal dataset [43], the polarimetrics thermal dataset extended to 111 subjects [11],[43]-ext. and a private dataset containing paired visible/thermal faces of 126 subjetcs. DAF was mainly compared to DPM [101] and shows significant improvement.
@@
7.3 Longwave with polarimetry
@@ 9 Open research challenges
Despite significant achievements, CFR still faces a number of challenges. In this section, we proceed to identify and discuss some pertinent challenges that are essential for further success of CFR.
@@
Scarcity of datasets. Unlike traditional FR, where millions of faces can be leveraged to train deep learning based models, CFR remains significantly impeded by lack of large-scale heterogeneous face datasets. In addition, very few existing datasets are endowed with annotations of face bounding box and landmarks [38]. Such annotations are essential in developing automated face and landmark detection for CFR.
@@
Large standoff distance. Limited work has focused on CFR at large standoff distances [126]. Most existing works have focused on CFR, with the assumption that faces have been captured at a short standoff distance. However, capturing faces at large standoff distances results in face image quality degradation, including low-resolution, blur and noise. Early attempts towards addressing this challenge have sought to restore high-quality images from degraded ones [126, 139]. Such methods were developed by learning a locally linear mapping between low-quality patches and high-quality patches. The linear mapping assumption has been relaxed on patch-level, while in practice, mapping is often considered to be non-linear. In this context, it has been demonstrated that GANs have been highly suitable to model non-linear mappings.
@@
Modality gap. The key problem with CFR is the apparent disparity between visible and invisible spectrum, and this strongly impacts the comparison performance (see Experiment 2). Although cross-spectrum featured-based methods, such as transfer learning or domain adaptation have achieved good results, especially in the Active IR band, such methods remain less efficient in the Passive IR band. An obvious solution is to incorporate details with regard to both geometrical and textural information. Unsupervised-processing methods with deep generative models, such as cross-spectrum images synthesis, appear to be the trend, in particular as a result of GAN-improvement. So far, ”synthetic data” has not yet outperformed the CFR state-of-the-art, but we foresee promising results.
@@
Privacy. Infrared imaging is a keystone biometric solution in terms of security and privacy [6]. Firstly, humans are not able to recognize someone as the IR-appearance does not resemble RGB-images (Figure 4). Secondly, it allows to ensure pseudo-anonymized recognition and compliance with legal requirements regarding personal data, as well as template protection. Finally, IR is robust to attempted diversion and spoofing attacks, as artifacts are directly identifiable by appearing differently on face images. While intuitive, the above has not been studied rigorously in scientific literature.
@@
Presentation Attacks. FR systems have shown to be vulnerable to presentation attacks, where an adversarial uses artifacts such as print, video replay, makeup and 3D mask to undermine the system. Reflection difference between live faces and spoof faces under NIR [140, 141], SWIR [142] or LWIR [8] spectral bands can be exploited to detect such presentation attacks. CFR can provide a unique insight, complementing RGB-imagery, as it imparts liveness detection, as well as shape determination [22].
@@
Next generation of synthesis models for CFR. As the field of deep learning continues to evolve, new and innovative methods for generating synthetic images are emerging. While GANs have been popular in recent years for translating thermal face images into visible-like face images [106, 120, 143, 10, 70], they still face several limitations such as mode collapse and instability during training. In contrast, models based on the diffusion process, such as the Diffusion Probabilistic Models (DPMs), offer a promising alternative for generating high-quality synthetic images with greater stability and less distortion (identity better preserved). Hence, transitioning from GANs to DPMs opens up exciting new possibilities for generating interspectral images with improved realism and accuracy, rendering them a worthwhile avenue of future research with new scientific keystones. In this context, a pioneer work has explored DPM proposing a Denoising Diffusion Probabilistic Model [144]. By exploring and developing these new models, the field of computer vision will advance, towards creating more sophisticated and reliable systems for a wide range of applications related to CFR.

### v4

@@ 1 Introduction
Face Recognition (FR) has been a very active research field for the last several decades [1, 2, 3, 4]. Advances in deep convolutional neural networks (CNNs) and the associated seminal work on DeepFace [5] have brought significant progress in this area, tackling a number of challenges including variations in pose, illumination and expression (PIE), as well as resolution and unconstrained settings. While such work have predominantly focused on the visible (VIS) spectrum, considering additional spectra allows for increased robustness [6, 7], in particular in the presence of different poses, illumination variations, noise, as well as occlusions. Further benefits include incorporating the absolute size of objects, as well as robustness to presentation attacks such as the use of makeup and masks to circumvent a system [8]. Therefore, comparing RGB face images against face images acquired beyond the visible spectrum, often referred to as Cross-spectral Face Recognition (CFR), is of particular significance in designing FR systems for defense, surveillance, and public safety [6] and falls under the broader category of Heterogeneous Face Recognition (HFR) that addresses the broader concept of recognizing faces across diverse and non-uniform conditions. Specifically, heterogeneous refers to face images being represented in different modalities, e.g., visible, infrared, 2D-3D, resolutions, hand-drawn sketches, and depth (see Figure 1).
The focus of this survey is on CFR, and specifically on comparing visible face images with images acquired beyond the visible spectrum, namely the infrared spectrum. We note that CFR is more challenging than traditional FR for both, human examiners, as well as computer vision algorithms due to the following three factors. Firstly, there is large intra-spectral variation, where within the same modality, face samples of the same subject can exhibit larger appearance variations than face samples of different subjects. Secondly, the modality gap is of concern, where appearance variation between two face samples of the same subject can be larger than that of two samples belonging to two different subjects (see Figure 1). This can result in degraded face recognition performance. Finally, the limited availability of training samples of cross-modality face image pairs can make it difficult to design effective CFR methods using deep neural networks.
Figure 1: Examples of heterogeneous face images of the same subject captured in different modalities, i.e., domains. Images are from the MCXFace dataset [9].
@@
Recent advances in CNNs and generative adversarial networks (GANs) have allowed for remarkable improvements in CFR [10, 11, 12, 13]. The goal of this survey article is to provide an overview of these advances. Specifically, we describe the proposed methods, used datasets, as well as the results and associated insights. In this context, we consider the electromagnetic spectrum ranging from near infrared (NIR) to more challenging bands such as midwave infrared (MWIR) and longwave infrared (LWIR), as illustrated in Figure 2. Unlike other introductory overview articles including [14, 6, 15], we focus on highlighting recent advances based on deep learning in addressing CFR, and techniques that have been developed to bridge the modality gap. Table 1 revisits existing survey papers and highlights novel contributions we present in this survey.
In particular, we showcase the two main approaches in CFR: (i) cross-spectrum feature-based algorithms, which extract and compare features, and (ii) cross-spectrum image synthesis algorithms, which generate images from one spectrum to another for matching and recognition. The former involves comparing infrared probes against visible face image galleries within a common feature subspace. However, following the success of GAN-based image synthesis, the latter attempts to synthesize a “pseudo-visible” images from infrared images. These synthesized face images allow both, academic and commercial off-the-shelf (COTS) FR systems trained on the visible spectrum to be used for comparison.
Table 1: Summary of other related overview articles and the novelty of our work.
Survey reference Description Novelty proposed by our survey
R.S. Ghiass et al. [14] (2014) Infrared Face Recognition: A Comprehensive Review of Methodologies and Databases Reviews methodologies and databases for infrared face recognition, discussing the robustness of thermal imaging to facial expressions, pose variations, and the use of blood vessel patterns. Incorporation of Recent Deep Learning Advances: We introduce the latest deep learning techniques, specifically tailored for CFR, surpassing traditional methods reviewed [14]. Focus on Multimodal Integration: We explore the integration of multiple spectral bands, aimed at enhancing recognition accuracy and robustness, a more comprehensive approach than prior reviews.
S. Ouyang et al. [16] (2016) A Survey on Heterogeneous Face Recognition: Sketch, Infra-red, 3D and Low-resolution Provides a detailed description of datasets and benchmarks relevant to evaluation and offers an extensive review of established techniques. Focus on CFR: Our survey is focused specifically on CFR including NIR, SWIR, MWIR, LWIR and Polar-LWIR, compared to NIR [16]. Deep Learning-based Approaches: We revisit recently proposed deep learning methods for CFR, as opposed to traditional pattern recognition methods [16].
S. Hu et al. [6] (2017) Heterogeneous Face Recognition: Recent Advances in Infrared-to-Visible Matching Focuses on methods for matching infrared images to visible spectrum images, highlighting recent advances and challenges in heterogeneous face recognition. Bridging the Modality Gap: We provide an in-depth analysis of state-of-the-art techniques for bridging the modality gap using deep learning. Comprehensive Dataset Review: We extensively revisit newly available datasets, enhancing the evaluation framework for CFR methods.
R. Munir and R.A. Khan [15] (2019) An Extensive Review on Spectral Imaging in Biometric Systems: Challenges and Advancements Discusses challenges and advancements in spectral imaging for biometric systems, including various spectral bands and their applications. Detailed Exploration of Cross-Spectral Techniques: We delve into CFR methods with a focus on practical applications. Future Directions and Challenges: We outline future research directions and unresolved challenges, providing a road map for ongoing and future work in the field. Comparison of Spectral Bands: We here compare the effectiveness of different spectral bands, providing insights into optimal spectrum utilization for face recognition.
@@
The remainder of this paper is organized as follows. We formalize CFR in Section 2 and present some applications in Section 3. We discuss the various spectral bands, i.e., sensing modalities, that are relevant to face recognition, in Section 4. In Section 5 we describe the general modules of a CFR system, viz., face image pre-processing and feature extraction. The various deep learning based CFR methods that have been proposed in the literature are explained in Section 6.1 and 6.2, while the relevant datasets used for research in this field are presented in Section 7. Finally, in Section 8, we discuss open research problems that are currently being addressed in the field of CFR.
@@ 3 Applications of CFR
Figure 3: A monitoring system equipped with a thermal sensor applies face and facial landmark detection to a video sequence captured in the wild at night, successfully detecting the person despite challenging conditions, paving the way for subsequent identification steps. Images are from Anghelone et al. [20].
FR-systems have been widely deployed and operate mainly in the visible spectrum [21], due to the ubiquity of low-cost cameras. As motivated in the introduction, CFR-systems have been beneficial in additional applications, for example, in Intelligence Video Surveillance (IVS). Beyond access control where acquisition is controlled w.r.t. user collaboration, the application of CFR has strong potential to provide new capabilities in law enforcement, military and the AI community, especially for commercial purposes. In this context, IR sensors offer several advantages over traditional visible light imaging. They enable long-range monitoring and capture biometric modalities at variable standoff distances. IR sensors can also capture images through certain types of obscuring materials and perform well in low-light or nighttime scenarios. However, identifying individuals in real-world situations often involves uncooperative behavior. The unconstrained environment necessitates capturing images of subjects at a distance from the camera, accounting for various facial poses, expressions, and even occlusions.
Law Enforcement and Homeland Security
A pertinent application for CFR has been in law enforcement and homeland security [22]. In this context, IR-sensors can be used for long-range monitoring and capturing biometric modalities at variable standoff distances, as well as in low-light and night-time scenarios. These settings are depicted in Figure 3. In such cases, the acquired image often has to be compared against a gallery consisting of visible spectrum face images for identification purposes. This capability enhances the effectiveness of monitoring suspects and ensuring security in public spaces.
Military Applications
Similar applications are found in military settings involving identifying individuals in challenging environments [7]. Night-vision devices, which often operate in the infrared spectrum, can capture images in complete darkness. These images can be cross-referenced with visible spectrum databases, facilitating the identification and tracking of persons of interest in combat or surveillance operations.
Commercial Applications
In the commercial sector, CFR can improve security measures in places such as shopping malls and airports. AI-powered security systems using CFR can manage access control more effectively, recognizing individuals across different lighting conditions and enhancing overall safety.
Presentation Attack Detection
CFR-systems are also beneficial for deflecting presentation attacks [23, 24], which involve attempts to deceive biometric systems using photographs, masks, or other fraudulent methods. By leveraging images from multiple spectral bands, these systems can better discern between genuine and fake biometric data, enhancing security measures against such attacks.
Sketch-Visual Identification
On a wider scale, HFR (which is a superset of CFR) offers notable application in sketch-visual identification. Law enforcement agencies can use HFR to match a suspect’s sketch with a visible spectrum face image from a mugshot database. This method is particularly useful when photographic evidence is unavailable, and only a hand-drawn sketch of the suspect exists [25].
@@ 5.3 Face Comparison
6 Algorithms of Different Spectra
@@
6.1 Reflective IR-to-Visible FR
6.1.1 Near Infrared
@@
IDNet. Reale et al. [47] utilized GoogLeNet with small convolutional filters, pre-trained on the visible CASIA WebFace dataset. For the NIR-to-VIS CFR scenario, they initialized two identical networks based on the pre-trained model, excluding the fully connected and softmax layers. The outputs from these two networks, termed VisNet and NIRNet, were concatenated to form a Siamese network with a contrastive loss function. This coupled deep network design effectively mapped NIR and VIS faces into a spectrum-independent feature space.
@@
CFC. He et al. [12] proposed Cross-spectral Face Completion (CFC), a GAN-based method, aimed at synthesizing visible spectrum (VIS) images from near-infrared (NIR) images. This heterogeneous face synthesis process comprises two complementary components: (a) texture inpainting, recovering VIS facial textures from NIR images, and (b) pose correction, mapping any pose in NIR images to a frontal pose in VIS images. These components are fused using a warping procedure within an adversarial network, regularized by UV loss, adversarial loss, and L1 reconstruction loss. The main contributions of this work include creating a novel GAN-based end-to-end framework for cross-spectral face synthesis without the need to assemble multiple image patches. The framework includes encoder-decoder structured generators and two novel discriminators to fully account for variations between NIR and VIS images. This approach simplifies the unsupervised heterogeneous face synthesis problem to a one-to-one image translation problem. The decomposition of texture inpainting and pose correction enables the generation of realistic, identity-preserving VIS face images.
@@ 6.1.2 Shortwave Infrared
VGGFace. Bihn et al. [53] explored the use of a pre-trained VGG-Face network on visible images to extract features from shortwave infrared (SWIR) images. They hypothesized that the VGG-Face network would perform well on both VIS and SWIR face images since their facial appearance differences are less significant compared to MWIR and LWIR images. SWIR wavelengths at 935 nm, 1060 nm, 1300 nm, and 1550 nm were evaluated, along with a composite image combining 1060 nm, 1300 nm, and 1550 nm. Deep features were extracted from the fc7 layer of the VGG-Face network, producing a 4096-dimensional feature vector. The Euclidean distance between feature representations of VIS and SWIR images was used for similarity comparison.
@@
6.2 Emissive IR-to-Visible FR
6.2.1 Midwave Infrared
Appearance difference between MWIR and VIS is more pronounced than NIR and SWIR (Figure 5). Therefore, learning a common feature subspace for MWIR-to-VIS is a challenging problem. In most cases, methods developed over the years have relied on GANs to synthesize visible face images from thermal face images. Illustrative examples are shown in Figure 14 and 16. We report the related MWIR-to-VIS face comparison performance in Table 7.
@@ 6.2.2 Longwave Infrared
[…]
@@
DAF. Fondje et al. [120] proposed a Domain Adaptation Framework (DAF) featuring a new feature mapping sub-network with: (a) a cross-domain identification loss for effective ”co-registration” and ”synchronization,” and (b) a domain invariance loss for cross-domain regularization. Feature representations within each domain were extracted using a truncated pre-trained CNN (VGG16 or ResNet50), aiming to preserve discriminative information across visible and thermal spectra. DAF was embedded in the Residual Spectral Transform (RST), a residual block that transforms features with three specific 1x1 convolutions to enhance discriminability. Experiments using three datasets (polarimetric thermal, extended polarimetric thermal with 111 subjects, and a private dataset with 126 paired visible/thermal faces) showed significant improvement compared to DPM [103].
@@
6.2.3 Longwave with polarimetry
@@ 8 Open Research Challenges
Despite significant achievements, CFR still faces a number of challenges. In this section, we proceed to identify and discuss some key challenges that are essential for further success of CFR.
@@
Scarcity of datasets. Unlike traditional FR, where millions of faces can be leveraged to train deep learning based models, CFR remains significantly impeded by lack of large-scale heterogeneous face datasets. In addition, very few existing datasets include annotations of face bounding box and landmarks [41]. Such annotations are essential for developing automated face and landmark detection for CFR.
@@
Large standoff distance. Limited work has focused on CFR at large standoff distances [128]. Most existing studies assume that faces are captured at short standoff distances. However, capturing faces at large standoff distances results in face image quality degradation, including low-resolution, blur and noise. Early attempts towards addressing this challenge have sought to restore high-quality images from degraded ones [128, 141]. Such methods were developed by learning a locally linear mapping between low-quality patches and high-quality patches. The linear mapping assumption has been relaxed at the patch-level, while in practice, the mapping is often considered non-linear. In this context, it has been demonstrated that GANs are highly suitable to model non-linear mappings.
@@
Modality gap. The key problem with CFR is the apparent disparity between visible and invisible spectrum, which strongly impacts the comparison performance (see Experiment 2). Although cross-spectrum featured-based methods, such as transfer learning or domain adaptation have achieved good results, especially in the Active IR band, they remain less efficient in the Passive IR band. An obvious solution is to incorporate details related to both geometrical and textural information. Unsupervised-processing methods with deep generative models, such as cross-spectrum images synthesis, appear to be a promising trend, particularly with improvements in GANs. So far, ”synthetic data” has not yet outperformed the state-of-the-art in CFR, but we foresee promising results in the near future.
@@
Privacy. Infrared imaging is a keystone biometric solution in terms of security and privacy [6]. Firstly, humans are unable to recognize individuals from IR images, as IR appearances do not resemble RGB images (Figure 5). Secondly, IR imaging ensures pseudo-anonymized recognition and compliance with legal requirements regarding personal data and template protection. Finally, IR is robust against attempted diversion and spoofing attacks, as artifacts are directly identifiable by appearing differently on face images. While intuitive, the above has not been studied rigorously in scientific literature.
@@
Presentation Attacks. FR systems have shown to be vulnerable to presentation attacks, where adversaries use artifacts such as prints, video replays, makeup, and 3D masks to undermine the system. Reflection difference between live faces and spoof faces under NIR [142, 143], SWIR [144] or LWIR [8] spectral bands can be exploited to detect such presentation attacks. CFR can provide a unique insight, complementing RGB-imagery, as it imparts liveness detection, as well as shape determination [24].
@@
Next generation of synthesis models for CFR. As the field of deep learning continues to evolve, new and innovative methods for generating synthetic images are emerging. While GANs have been popular in recent years for translating thermal face images into visible-like face images [108, 122, 145, 10, 72], they still face several limitations such as mode collapse and instability during training. In contrast, models based on the diffusion process, such as the Diffusion Probabilistic Models (DPMs), offer a promising alternative for generating high-quality synthetic images with greater stability and less distortion, preserving identity better. Hence, transitioning from GANs to DPMs opens up exciting new possibilities for generating interspectral images with improved realism and accuracy, rendering them a worthwhile avenue of future research with new scientific keystones. In this context, a pioneer work has explored thermal to visible face translation by proposing a Denoising Diffusion Probabilistic Model [146]. By exploring and developing these new models, the field of computer vision will advance, towards creating more sophisticated and reliable systems for a wide range of applications related to CFR.